OpenStudy (anonymous):

Please help me!!!! Find the area under the curve y= 3/x on the interval [1,2]

OpenStudy (campbell_st):

you are being asked to integrate a log function so its \[A = \int\limits_{1}^{2} \frac{3}{x} ..dx\] which can be written as \[A = 3 \times \int\limits_{1}^{2} \frac{1}{x}..dx\] hope it helps

OpenStudy (sidsiddhartha):

A = ∫3/x dx-----------limits 1 to 2

OpenStudy (anonymous):

Okay so I just have to integrate that problem and I will have the answer???

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

awesome just one minute I am going to integrate it.

OpenStudy (sidsiddhartha):

so what's ur answer?

OpenStudy (anonymous):

Okay I got... \[\int\limits_{1}^{2} 1/x dx = \ln(2) =3\ln(2)\]

OpenStudy (anonymous):

I don't know if I did that right.

OpenStudy (sidsiddhartha):